# Evaluating Step-by-step Reasoning Traces: A Survey

### v2

@@ Abstract
Step-by-step reasoning is widely used to enhance the reasoning ability of large language models (LLMs) in complex problems. Evaluating the quality of reasoning traces is crucial for understanding and improving LLM reasoning. However, existing evaluation practices are highly inconsistent, resulting in fragmented progress across evaluator design and benchmark development. To address this gap, this survey provides a comprehensive overview of step-by-step reasoning evaluation, proposing a taxonomy of evaluation criteria with four top-level categories (factuality, validity, coherence, and utility). Based on the taxonomy, we review different evaluator implementations and recent findings, leading to promising directions for future research.
@@ 1 Introduction
  • Surveying existing evaluators for step-by-step reasoning evaluation based on their implementations, across diverse reasoning tasks and criteria (§4).
  • Identifying recent findings and promising directions for reasoning trace evaluation (§56).
@@ 2.1 Step-by-step reasoning
Step-by-step reasoning is where LLMs generate a series of intermediate natural language steps ("thoughts") before outputting the final answer (Wei et al., 2022). Each instance consists of two parts: a query and a reasoning trace, and the final answer. Upon seeing the query (user input), the LLM autoregressively generates its solution as a reasoning trace. Finally, a trace should include a final answer for the query, which can be compared to the ground truth. See Appendix A for details on different reasoning tasks.
@@ 3.3 Coherence
Coherence measures if a reasoning step’s preconditions are satisfied by the previous steps (Wang et al., 2023a; Lee and Hwang, 2025). For instance, if a trace includes the reasoning step "Next, we add 42 to 16." but the origin of the value 42 was never explained in the previous steps, this step is considered incoherent. An intuitive way to obtain an incoherent trace is randomly shuffling a coherent trace (Wang et al., 2023a; Nguyen et al., 2024), as the premise of some steps will not appear anywhere in the previous steps (incoherent) even though it can be eventually deduced from the query (valid).
Note that coherence judgment is inherently subjective and pragmatic compared to other criteria (Jacovi et al., 2024). For instance, seemingly trivial steps like "A part of something is present in that something" in WorldTree V2 (Xie et al., 2020) is annotated as necessary in Dalvi et al. (2021) but not necessary in Ott et al. (2023).
@@ 3.4 Utility
The narrower interpretation of utility is progress, or whether the step is correctly following the ground truth solution (Saparov and He, 2023; Nguyen et al., 2024). For instance, in Game of 24 (making the number 24 using 4 natural numbers and basic arithmetic operations) (Yao et al., 2023), a solution can be defined as a sequence of operations (e.g. $5+7=12$rightarrow $12-6=6$rightarrow $6\times 4=24$). In this task, the utility of a step (making $5+7=12$ from $5$ and $7$) can be directly assessed by checking if it is a part of a correct solution.
@@ 4 Evaluator types
Table 2: Evaluators for step-by-step reasoning, grouped by implementation type (E denotes that the method is an ensemble of different methods). The Domain column specifies the domains used for meta-evaluating each evaluator. "Arith" refers to arithmetic reasoning tasks, "Common" denotes commonsense question answering, and "Expert" corresponds to specialized expert domains like biomedical, legal, and financial reasoning (Appendix A). When evaluators are tested on both arithmetic and general math tasks, only "Math" is listed. The acronym FVCU in the Criteria column represents factuality, validity, coherence, and utility, respectively. For AutoRace (Hao et al., 2024), LLMs are instructed to list the criteria based on incorrect traces.
@@ 4.2 Intrinsic metrics
Uncertainty of the model can be used as an intrinsic proxy about the generated content’s quality (Xiao and Wang, 2021; Zhang et al., 2023b). Qiu et al. (2024) use token probability entropy, defined as $\Sigma_{t\in V}p(t)\textrm{log}(p(t))$ where $p$ is the probability distribution of all tokens in vocabulary $V$. Farquhar et al. (2024) and Kossen et al. (2024) extend the approach by clustering semantically similar tokens and calculating the entropy for each cluster. While uncertainty-based evaluators are primarily used for factuality (Wu et al., 2024a; Farquhar et al., 2024), they have also been applied for evaluating validity (Zhu et al., 2025) or utility (Hu et al., 2024), indicating that uncertainty can be a criteria-agnostic proxy of the quality of reasoning steps.
@@ 4.3 External evaluators
Cross-encoders simultaneously encode multiple sentences using a single, small network typically with millions of parameters (Devlin et al., 2019; Liu et al., 2019). They have been widely applied to solve tasks such as natural language inference (Bowman et al., 2015) and fact verification (Thorne et al., 2018), where one has to determine if the hypothesis can be inferred from the given premise. Cross-encoders trained on off-the-shelf tasks (Golovneva et al., 2023a; Zha et al., 2023; Prasad et al., 2023) or LLM-perturbed data (Zhu et al., 2024b) can be used to evaluate a reasoning step based on the query (groundedness) or previous steps (validity). However, their limited language understanding ability and shorter context length restrict their performance in more complex tasks.
@@ 5 Further improving evaluators
This paradigm can be extended to critic models that reason on reasoning traces, especially in meta-evaluation benchmarks. When applying majority voting of independently sampled $K$ evaluator scores in generative models (critic model, generative verifiers), the accuracy in predicting incorrect steps increases linearly with the scale of $\log K$ (Singhi et al., 2025; Kim et al., 2025b; She et al., 2025) (Figure 4, Majority voting). Furthermore, using large reasoning models (LRMs) with stronger reasoning capability by generating longer traces (Zheng et al., 2024a; Kim et al., 2025b; Khalifa et al., 2025) leads to significant improvement in error detection (Figure 4, Fine-tuned LRMs).
@@ 6 Future directions
Objective metrics for coherence evaluation.
LLMs often omit trivial inference steps in their reasoning (Saparov and He, 2023; Prasad et al., 2023), but even human annotators significantly disagree on whether a certain step can be taken as granted (Jacovi et al., 2024). As a result, we do not have many coherence evaluators or datasets (Table 1, 3). A human-annotated dataset of omittable and non-omittable steps would facilitate the development of practical coherence evaluators and partial context extractors.
@@ A.3 Others
Appendix B Meta-evaluation Datasets
@@
Table 3: List of PRM training data and meta-evaluation benchmarks with step-wise annotation. symbol indicates that the datasets include other tasks, such as summarization, instruction following, etc, where the # Trace column only counts the reasoning subset. Train/Eval columns denote if the dataset is used for training or meta-evaluation. Domain indicates what tasks are used to sample the reasoning trace. Criteria column shows the criteria used to annotate the data classified according to Section 3, where FVCU stands for factuality, validity, coherence, and utility, respectively. BiGGenBench (Kim et al., 2025a) applies query-specific evaluation criteria. Parentheses indicate that the criteria group is not explicitly distinguished in the labels. Human column indicates human annotation, where $\bullet$
@@
Datasets that annotate LLM-generated reasoning traces can be used to train an evaluator or perform meta-evaluation on different metrics. A summary of existing datasets can be found in Table 3. This section mainly focuses on the dataset construction process, namely data annotation and label types.
@@ B.1.2 Inter-annotator agreement
In many cases, the disagreement is significantly affected by the linguistic aspects of the reasoning trace. REVEAL (Jacovi et al., 2024) manually classifies steps that aroused disagreement between annotators into 13 distinct categories. Among these, frequent disagreement types like "World knowledge (some world knowledge might not be taken for granted)" and "Unclear reference (one proper noun can refer to multiple real-world entities)" are typical disagreement types observed in simpler recognizing textual entailment (natural language inference) tasks (Camburu et al., 2018; Lee et al., 2025), showing that these vagueness is present even in minimal settings.
@@ C.2 Comparison to other definitions
Appendix D Details for Discussions
This section provides further details regarding the Discussions (Section 5), specifically Figure 4.
@@ D.2 Data source
  • Fine-tuned PRM scores are from She et al. (2025).

### v3

@@ Abstract
Step-by-step reasoning is widely used to enhance the reasoning ability of large language models (LLMs) in complex problems. Evaluating the quality of reasoning traces is crucial for understanding and improving LLM reasoning. However, existing evaluation practices are highly inconsistent, resulting in fragmented progress across evaluator design and benchmark development. To address this gap, this survey provides a comprehensive overview of step-by-step reasoning evaluation, proposing a taxonomy of evaluation criteria with four top-level categories (factuality, validity, coherence, and utility). Based on the taxonomy, we review different datasets, evaluator implementations, and recent findings, leading to promising directions for future research.
@@ 1 Introduction
  • Surveying existing datasets and evaluators for step-by-step reasoning evaluation based on their implementations, across diverse reasoning tasks and criteria (§45).
  • Identifying recent findings and promising directions for trace evaluation (§67).
@@ 2.1 Step-by-step reasoning
Step-by-step reasoning is where LLMs generate a series of intermediate natural language steps ("thoughts") before outputting the final answer (Wei et al., 2022). Each instance consists of two parts: a query and a reasoning trace, and the final answer as a part of the reasoning trace. Upon seeing the query (user input), the LLM autoregressively generates its solution as a reasoning trace. Finally, a trace should include a final answer for the query, which can be compared to the ground truth. See Appendix A for details on different reasoning tasks.
@@ 3.2 Validity
Dataset Train Eval Domain Criteria # Trace Human
RAGTruth (Niu et al., 2024) $\bullet$ $\bullet$ Fact F 5.9k $\bullet$
HaluEval(Li et al., 2023a) $\bullet$ $\bullet$ Fact F 10k \ $\blacktriangle$
Math-Shepherd (Wang et al., 2024c) $\bullet$ Math U 440k $\times$
PRM800k (Lightman et al., 2024) $\bullet$ $\bullet$ Math V 75k $\bullet$
REVEAL (Jacovi et al., 2024) $\bullet$ Common FVC 3.4k $\bullet$
MATH-Minos (Gao et al., 2024a) $\bullet$ Math V 440k $\times$
SCDPO (Lu et al., 2024) $\bullet$ Math U 30k $\times$
MR-GSM8k (Zeng et al., 2024a) $\bullet$ Math V 3.0k $\bullet$
ProcessBench (Zheng et al., 2024a) $\bullet$ Math V 3.4k $\bullet$
MR-MATH (Xia et al., 2025) $\bullet$ Math VU 0.1k $\bullet$
PRMBench (Song et al., 2025) $\bullet$ Math VCU 6.2k \ $\blacktriangle$
VersaPRM (Zeng et al., 2025) $\bullet$ Expert FV 84.1k $\times$
Table 1: List of evaluator training data and meta-evaluation benchmarks. symbol indicates that the datasets include other tasks, such as summarization, instruction following, etc, where the # Trace column only counts the reasoning subset. Train/Eval columns denote if the dataset is used for training or meta-evaluation. Domain indicates what tasks are used to sample the reasoning trace. Criteria column shows the criteria used to annotate the data classified according to Section 3, where FVCU stands for factuality, validity, coherence, and utility, respectively. BiGGenBench (Kim et al., 2025a) applies hand-written, query-specific evaluation criteria (Custom). Human column indicates human annotation, where $\bullet$
\ $\blacktriangle$
@@ 3.3 Coherence
Coherence measures if a reasoning step’s preconditions are satisfied by the previous steps (Wang et al., 2023a; Lee and Hwang, 2025). For instance, if a trace includes the reasoning step "Next, we add 42 to 16." but the origin of the value 42 was never explained in the previous steps; this step is considered incoherent. An intuitive way to obtain an incoherent trace is randomly shuffling a coherent trace (Wang et al., 2023a; Nguyen et al., 2024), as the premise of some steps will not appear anywhere in the previous steps (incoherent) even though it can be eventually deduced from the query (valid).
Note that coherence judgment is inherently subjective and pragmatic compared to other criteria (Jacovi et al., 2024). For instance, seemingly trivial steps like "A part of something is present in that something" in WorldTree V2 (Xie et al., 2020) are annotated as necessary in Dalvi et al. (2021) but not necessary in Ott et al. (2023).
@@ 3.4 Utility
The narrower interpretation of utility is progress, or whether the step is correctly following the ground truth solution (Saparov and He, 2023; Nguyen et al., 2024). For instance, in Game of 24 (making the number 24 using 4 natural numbers and basic arithmetic operations) (Yao et al., 2023), a solution can be defined as a sequence of operations (e.g. $5+7=12$ $\rightarrow$ 126=612-6=6 $\rightarrow$ $6\times 4=24$). In this task, the utility of a step (making $5+7=12$ from $5$ and $7$) can be directly assessed by checking if it is a part of a correct solution.
@@
4 Meta-evaluation Datasets
Datasets that annotate LLM-generated reasoning traces serve as key resources for training evaluators and conducting meta-evaluations between evaluators. A summary of existing datasets is provided in Table 4.
Among these, one of the most influential is PRM800k (Lightman et al., 2024). PRM800k consists of crowdsourced tertiary validity labels (positive, negative, neutral) assigned step by step, framing reasoning trace error detection as a sequence classification problem. Its design has inspired several successors (Zeng et al., 2024a; Xia et al., 2025), setting the paradigm for subsequent reasoning trace evaluation resources.
To address different needs, several extensions have been developed. Since human annotations are costly and difficult to scale, many works have explored automatic labeling—either by estimating utility through Monte Carlo Tree Search (MCTS) (Wang et al., 2024c; Luo et al., 2024b; Setlur et al., 2024) or by generating perturbed traces with LLMs (Lu et al., 2024; Song et al., 2025). More recent datasets further broaden the scope by enabling multi-criteria meta-evaluation (Jacovi et al., 2024; Tyen et al., 2024; Song et al., 2025) and expanding coverage beyond mathematics into diverse domains (Zeng et al., 2024b, 2025).
Additional details are provided in Appendix B.
@@ 5 Evaluator types
Table 3: Evaluators for step-by-step reasoning, grouped by implementation type (E denotes that the method is an ensemble of different methods). The Domain column specifies the domains used for meta-evaluating each evaluator. "Arith" refers to arithmetic reasoning tasks, "Common" denotes commonsense question answering, and "Expert" corresponds to specialized expert domains like biomedical, legal, and financial reasoning (Appendix A). When evaluators are tested on both arithmetic and general math tasks, only "Math" is listed. The acronym FVCU in the Criteria column represents factuality, validity, coherence, and utility, respectively. For AutoRace (Hao et al., 2024), LLMs are instructed to list the criteria based on incorrect traces (Custom).
@@ 5.2 Intrinsic metrics
Uncertainty of the model can be used as an intrinsic proxy for the generated content’s quality (Xiao and Wang, 2021; Zhang et al., 2023b). Qiu et al. (2024) use token probability entropy, defined as $\Sigma_{t\in V}p(t)\textrm{log}(p(t))$ where $p$ is the probability distribution of all tokens in vocabulary $V$. Farquhar et al. (2024) and Kossen et al. (2024) extend the approach by clustering semantically similar tokens and calculating the entropy for each cluster. While uncertainty-based evaluators are primarily used for factuality (Wu et al., 2024a; Farquhar et al., 2024), they have also been applied for evaluating validity (Zhu et al., 2025) or utility (Hu et al., 2024), indicating that uncertainty can be a criteria-agnostic proxy of the quality of steps.
@@ 5.3 External evaluators
Cross-encoders simultaneously encode multiple sentences using a single, small network often with millions of parameters (Devlin et al., 2019; Liu et al., 2019). They have been widely applied to solve tasks such as natural language inference (Bowman et al., 2015) and fact verification (Thorne et al., 2018), where one has to determine if the hypothesis can be inferred from the given premise. Cross-encoders trained on off-the-shelf tasks (Golovneva et al., 2023a; Zha et al., 2023; Prasad et al., 2023) or LLM-perturbed data (Zhu et al., 2024b) can be used to evaluate a reasoning step based on the query (factuality) or previous steps (validity). However, their limited language understanding ability and shorter context length restrict their performance in more complex tasks.
@@ 6 Further improving evaluators
This paradigm can be extended to critic models that reason on reasoning traces, especially in meta-evaluation benchmarks. When applying majority voting of independently sampled $K$ evaluator scores in generative models (critic model, generative verifiers), the accuracy in predicting incorrect steps increases linearly with the scale of $\log K$ (Singhi et al., 2025; Kim et al., 2025b; She et al., 2025) (Figure 4, Majority voting). Furthermore, using large reasoning models (LRMs) with stronger reasoning capability by generating longer traces (Zheng et al., 2024a; Kim et al., 2025b; Khalifa et al., 2025) leads to significant improvement in error detection (Figure 4, Scaling (LRMs)).
@@ 7 Future directions
Advanced methods for finding premises.
NLI-based validity and coherence evaluation significantly benefit from determining the previous steps that the current step uses as a premise (Mukherjee et al., 2025). However, finding such steps is not a trivial task. ROSCOE (Golovneva et al., 2023a) uses the minimum NLI score of all (previous step, current step) combinations, which ignores cases where a step has multiple premises. Recent works (Ling et al., 2023; Tyen et al., 2024; Mukherjee et al., 2025) make the reasoner LLM annotate the premises of the given step. Plausible but underexplored approaches include applying uncertainty-based methods (Chen et al., 2023; Wu et al., 2024a) or training a parser that annotates the logical dependencies between steps as graphs (Lee et al., 2025b).
@@
Rubric-based evaluation for complex and expert-level tasks.
Existing evaluators often apply identical evaluations for all reasoning trace, e.g., using the same LLM-as-a-judge prompt for all inputs. However, as the reasoning tasks require more domain knowledge and expertise, there is an increasing need for highly specific rubrics for evaluating reasoning traces (Kim et al., 2025a). For instance, one can calculate the sum of an arithmetic sequence by adding all terms one by one or finding a general term; the problem-specific rubrics explicitly prefer the latter. However, manual rubric generation is costly and less scalable, which motivates automatic extraction/generation of high-quality reasoning trace rubrics. AutoRace (Hao et al., 2024) aims to generate rubrics automatically based on incorrect responses, while RaR (Gunjal et al., 2025) extracts checklist-style rubrics from ground-truth biomedical documents. Still, automatically obtaining expert-level, high-quality rubrics for more diverse reasoning tasks remains an open question.
@@ A.3 Others
Appendix B Appendix for Meta-evaluation Datasets
@@
Table 4: List of evaluator training data and meta-evaluation benchmarks. symbol indicates that the datasets include other tasks, such as summarization, instruction following, etc, where the # Trace column only counts the reasoning subset. Train/Eval columns denote if the dataset is used for training or meta-evaluation. Domain indicates what tasks are used to sample the reasoning trace. Criteria column shows the criteria used to annotate the data classified according to Section 3, where FVCU stands for factuality, validity, coherence, and utility, respectively. BiGGenBench (Kim et al., 2025a) applies hand-written, query-specific evaluation criteria (Custom). Human column indicates human annotation, where $\bullet$
@@
This appendix includes discussions on the dataset construction process, with a focus on data annotation and label types. A summary of existing datasets can be found in Table 4.
@@ B.1.2 Inter-annotator agreement
In many cases, the disagreement is significantly affected by the linguistic aspects of the reasoning trace. REVEAL (Jacovi et al., 2024) manually classifies steps that aroused disagreement between annotators into 13 distinct categories. Among these, frequent disagreement types like "World knowledge (some world knowledge might not be taken for granted)" and "Unclear reference (one proper noun can refer to multiple real-world entities)" are typical disagreement types observed in simpler recognizing textual entailment (natural language inference) tasks (Camburu et al., 2018; Lee et al., 2025a), showing that these vagueness is present even in minimal settings.
@@ C.2 Comparison to other definitions
Appendix D Details for Section 6
This section provides further details regarding the Section 6, specifically Figure 4.
@@ D.2 Data source
  • Fine-tuned Critic Model scores are from She et al. (2025).